# How much can we forget about Data Contamination?

### v1

@@ 5.1 Weight Decay as a Mechanism for Forgetting Training Examples
We now analyze the weight-decay mechanism of forgetting in different LLM training runs. Of course, the decay of past gradients described in Proposition 1 is only one effect that contributes to forgetting. Indeed, there is also the potential for different gradient updates in the sum (3) to cancel each other. However, because a term decayed to zero is definitively forgotten, we can consider the cumulative weight decay as an upper bound on forgetting. Figure 5 depicts the evolution of the forgetting term $w_{t_{1}}^{t_{2}}$ for the 124M parameter model from Section 4.2, OLMo 7B, and Llama 3 405B (where we assume the model trained with a weight decay of 0.1). For the training data at each decline of the training run, Figure 5 depicts forgetting curves that indicate how much the corresponding gradients decay as we continue training. For the 124M model depicted in Figure 5(a), even the initialization is not completely decayed towards the end (the blue curve is still strictly larger than zero). In contrast, for OLMo-7B, the gradients of the first 40% of the training data decay to zero until the end of training, meaning that this data is forgotten (Figure 5(b)). Llama 3 405B also experiences significant decay of the early gradients (Figure 5(c)).
The interplay between the weight decay parameter and learning rate schedule of AdamW creates an interesting dynamic. Lower learning rates towards the end of training (1) slow down the forgetting of past training examples and (2) decrease the impact of later training examples on the final model weights. To better understand this dynamic, we plot the sum of the terms $\lambda_{t_{1}}w_{t_{1}}^{t_{2}}$ for each decile of the training run, normalized by the same sum over the entire training run. This is depicted in the bottom row of Figure 5 and can be thought of as a simple approximation of how the final model weights are composed by the gradients of different training deciles. Interestingly, this approximation suggests that the Llama 3 405B training run, where supposedly a lot of expertise has gone into the choice of the hyperparameters, results in a model where the approximate influence of different training steps is symmetrically distributed around the middle of training (Figure 5(c)). In contrast, the OLMo-7B training could seemingly benefit from further decaying the learning rate to give more weight to early versus late gradients.
While our analysis in this Section considers the mechanistic effect of individually decaying gradient updates, it does not model any interactions between different gradient updates. For example, if the weight updates at a later time step $t_{2}$ were aligned with past updates at $t_{1}$, then the model might not forget the information even if the effect of past updates vanishes from the sum. However, such complex interactions are avoided if the model updates from contaminated samples are orthogonal. Formalizing this observation leads to a more rigorous version of Proposition 1, presented in Supplement A.5. The argument is that under suitable gradient orthogonality conditions, the decay of past gradients can guarantee forgetting.
5.2 Practical forgetting occurs faster than what the theory predicts
(a) Empirical
(b) Theory
(c) Empirical & Theory
(d) Empirical & Theory
Figure 6: The theoretical estimates provide an upper bound on empirical forgetting, which can occur much faster. (a) Empirical forgetting for three different weight decay parameters. (b) The corresponding cumulative weight decay. (c) Empirical forgetting and theoretical estimate for the default weight decay of 0.1. The y-axis of the cumulative weight decay $w_{t_{1}}^{t_{2}}$ is calibrated to start at the peak of the empirically observed overfitting. (d) Empirical forgetting and theoretical estimate for OLMo-1B (HellaSwag).
In the previous Section 5.1, we have derived a simple theory of forgetting via cumulative weight decay. We now investigate how the theoretical estimates relate to the empirically observed forgetting.
The main parameter that controls the theoretical forgetting curves is the weight decay parameter. Therefore, we ask how forgetting changes empirically when we change the weight decay. Figure 6(a) depicts the result of repeating the forgetting experiment from Section 4.2 with three different choices for the weight decay parameter. From Figure 6(a), we see that the weight decay parameter controls the impact of contamination at all time steps, where a larger weight decay parameter leads to more forgetting and a smaller weight decay parameter to less forgetting. This is consistent with the theoretical predictions depicted in Figure 6(b), meaning there is a qualitative alignment between the empirical results and our theoretical predictions.
To better understand the quantitative relation between empirical forgetting and the theoretical estimates, we ask how the empirically forgotten fraction (of cross-entropy loss or accuracy) relates to the cumulative weight decay. Figure 6(c) depicts the empirical decay and corresponding theoretical prediction for the model from Section 4.2. We see that the theoretical estimate is somewhat pessimistic and that forgetting occurs faster than what is predicted by the theory. Figure 6(d) is similar to Figure 6(c), except that we consider the OLMo-1B forgetting experiment from Section 4.3. Figure 6(d) depicts a case where forgetting occurs much faster than what is predicted by the theory. Interestingly, we also see that the empirical rate of forgetting, at least in this experiment, is not smaller for the larger model.
@@ 6 Discussion
This work presents different experiments on data contamination and forgetting. We have seen that the impact of contamination can vanish as the size of the training data increases – an aspect that has largely been overlooked in the literature [71, 57, 40]. We have also shown that the hyperparameters of AdamW play an important part in forgetting – an insight that might inform the parametrization of future training runs. Of course, it would be interesting to study contamination and forgetting on a larger scale. Nevertheless, we have shown that the effects of contamination can vanish even though the number of steps we could train the models for was limited.
We have studied data contamination with a focus on the leakage of benchmark questions into the training data. This means that our work might be more informative about the topic than other works that study contamination in different contexts. At the same time, one has to be careful when extrapolating our results, especially to a privacy setup [12, 39]. This is because empirical forgetting might behave differently for random strings or otherwise uniquely identifiable information [13].
The “contamination problem” as studied in this paper has interesting connections to many areas of machine learning, including privacy [29, 39], data attribution [41], and generalization [9, 33, 53]. The connection to data attribution is especially relevant in light of our results. This is because we demonstrate cases where the presence or absence of a datapoint in the training data is irrelevant for the model behavior on that same datapoint, meaning it does not make sense to attribute model behavior to individual datapoints in this regime.
@@ A.3 Filtering Near-Duplicate Benchmark Questions
A.4 Proof of Proposition 1
Proposition 2.
(The Decay of Past Gradients) The number of optimization steps $T=t_{2}-t_{1}$ that are required to make the contribution of a model update at time $t_{1}$ small, that is $w^{t_{2}}_{t_{1}}\leq\epsilon$ for some small ϵ+italic-$\epsilon\in\mathbb{R}^{+}$, scales as $T\gtrsim\frac{log(1/\epsilon)}{\gamma\lambda_{\text{avg}}}$, where $\lambda_{\text{avg}}=\frac{1}{T}\sum_{t={t_{1}}}^{t_{2}}\lambda_{t}$ is the average learning rate of the optimizer between $t_{1}$ and $t_{2}$.
@@ A.4 Proof of Proposition 1
Without loss of generality, mapping $t_{1}=1$ and $T=t_{2}$, we have from equation 3:
$w^{T}_{1}=\prod_{i=1}^{T}(1-\lambda_{i}\gamma)$
Assigning the forgetting ratio to be less than $\epsilon$ according to our criteria, we have:
$\prod_{i=1}^{T}(1-\lambda_{i}\gamma)$ $\leq\epsilon$
$\sum_{i=1}^{T}\log(1-\lambda_{i}\gamma)$ $\leq\log\epsilon$
$\sum_{i=1}^{T}(-\lambda_{i}\gamma)$ $\lesssim\log\epsilon~{}~{}~{}~{}(\log(1-x)\approx-x\text{~{}for small~{}}x)$
$T\times\left(\frac{1}{T}\sum_{i=1}^{T}\lambda_{i}\right)\times\gamma$ $\gtrsim\log\frac{1}{\epsilon}$
Re-arranging this equation gives us the desired result, where $\lambda_{avg}=(\frac{1}{T}\sum_{i}\lambda_{i})$. ∎
A.5 Extended Analysis of Forgetting & Gradient Alignment
To understand the effect of weight decay on forgetting, we analyze two different stages of optimization: (1) the contamination stage, where the training set consists of only the contaminated samples, and (2) the forgetting stage, where the training set consists only of clean samples. Here, we consider the SGD learning algorithm, but the resulting analysis also applies to SGD with momentum. In particular, to illustrate the effect of weight decay, we assume the usage of SGD for the contamination stage and SGD with weight decay for the forgetting stage.
We now introduce some notation. Let $\theta\in\mathbb{R}^{D}$ be the weights of the model, and let $\mathcal{X}_{\text{cont}}=\{\mathbf{x}^{\text{cont}}_{1},\mathbf{x}^{\text{cont}}_{2},...\mathbf{x}^{\text{cont}}_{N_{\text{cont}}}\}$ be the contamination set and $\mathcal{X}_{\text{clean}}=\{\mathbf{x}^{\text{clean}}_{1},\mathbf{x}^{\text{clean}}_{2},...\mathbf{x}^{\text{clean}}_{N_{\text{clean}}}\}$ be the clean pre-training set. The training data used for the contamination stage is thus $\mathcal{X}_{\text{cont}}$, and the training data for the forgetting stage is $\mathcal{X}_{\text{clean}}$. Let $\ell(\mathbf{x}_{i})\in\mathbb{R}$ be the loss associated with sample $\mathbf{x}_{i}$. Let the model be initialized at the contamination stage with $\theta=\theta_{\text{init}}$. The learning algorithm is (single batch size) SGD, which is run for a total of $N_{\text{cont}}$ steps for the contamination stage and $N_{\text{clean}}$ steps for the forgetting stage. First, we observe that the weights at the end of the contamination stage are:
$\theta^{\prime}=\theta^{\text{init}}-\underbrace{\sum_{i=1}^{N_{cont}}\lambda_{i}{\nabla_{\theta}}_{i}\ell(\mathbf{x}^{\text{cont}}_{i})}_{\theta^{\text{cont}}}$
We thus denote the weights at the end of the contamination stage as $\theta^{\prime}=\theta^{\text{init}}+\theta^{\text{cont}}$. For the subsequent fine-tuning stage to forget information regarding samples $\mathcal{X}_{\text{cont}}$, we first define a criterion to identify forgetting based on the angle between a weight vector and the contaminated part identified above.
Forgetting Criteria. A model with weights $\theta$ is said to have "forgetten" information contained in $\theta^{\text{cont}}$ if $\frac{|\theta^{\top}\theta^{\text{cont}}|}{\|\theta^{\text{cont}}\|^{2}_{2}}\leq\epsilon$, which we call the "forgetting ratio". Here, ϵ+italic-$\epsilon\in\mathbb{R}^{+}$ is a small constant.
We now proceed with an analysis of the forgetting stage. To enable this, we make the following important assumption that the gradients of clean and contaminated samples are orthogonal for all clean and contaminated samples across all optimization steps. This is a relatively strong assumption, and it quantifies the intuition that model updates required by SGD to memorize clean samples and contaminated samples are distinct.
[…]
We also make another minor simplifying assumption that the weight initialization is orthogonal to both these quantities.
[…]
Given these assumptions, we are ready to state our result.
Proposition 3.
(Forgetting Time) The number of optimization steps $T_{\text{forget}}$ in the forgetting stage, such that the weights $\theta_{T_{\text{forget}}}$ satisfy the $\epsilon$-forgetting criteria is given by: $T_{\text{forget}}\gtrsim\frac{\log(1/\epsilon)}{\lambda_{\text{avg}}\gamma}$, where $\lambda_{\text{avg}}=\frac{1}{T}\sum_{i=1}^{T}\lambda_{i}$ is the average learning rate of the optimizer.
Proof.
We first compute the forgetting ratio at $\theta=\theta^{\prime}$, and as a consequence of Assumption 2, verify that the forgetting ratio is equal to one.
Let us now denote these weights as $\theta^{\prime}_{0}=\theta^{\prime}$, used as initialization for the forgetting stage. Analyzing the first optimization step, and the subsequent forgetting ratio, we have:
[…]
$=\frac{|({(\theta^{\text{init}}+\theta^{\text{cont}})-\lambda_{0}{\nabla_{\theta}}_{0}\ell(\mathbf{x}^{\text{clean}}_{0})-\lambda_{0}\gamma((\theta^{\text{init}}+\theta^{\text{cont}}))})^{\top}\theta^{\text{cont}}|}{\|\theta^{\text{cont}}\|^{2}_{2}}$
$=(1-\lambda_{0}\gamma)~{}~{}~{}~{}~{}~{}(\text{From Assumptions 1\& 2})$
We can similarly analyze the subsequent optimization steps to compute the forgetting ratio, which for some step $t+1$ is:
$\underbrace{\frac{|{\theta^{\prime}_{t+1}}^{\top}\theta^{\text{cont}}|}{\|{\theta^{\text{cont}}}\|^{2}_{2}}}_{\text{Forgetting ratio at $t+1$}}$ $=\frac{|({\theta^{\prime}_{t}-\lambda_{t}{\nabla_{\theta}}_{t}\ell(\mathbf{x}^{\text{clean}}_{t})-\lambda_{t}\gamma\theta^{\prime}_{t}})^{\top}\theta^{\text{cont}}|}{\|\theta^{\text{cont}}\|^{2}_{2}}$
$=\underbrace{\frac{|{\theta^{\prime}_{t}}^{\top}\theta^{\text{cont}}|}{\|{\theta^{\text{cont}}}\|^{2}_{2}}}_{\text{Forgetting ratio at $t$}}(1-\lambda_{t}\gamma)$
Unrolling the recurrence till step $T$, we have that:
${\frac{|{\theta^{\prime}_{T}}^{\top}\theta^{\text{cont}}|}{\|{\theta^{\text{cont}}}\|^{2}_{2}}}=\underbrace{\frac{|{\theta^{\prime}_{0}}^{\top}\theta^{\text{cont}}|}{\|{\theta^{\text{cont}}}\|^{2}_{2}}}_{\text{= 1}}\times\prod_{i=1}^{T}(1-\lambda_{i}\gamma)$
Assigning the forgetting ratio to be less than $\epsilon$ according to our criteria, we have:
$\prod_{i=1}^{T}(1-\lambda_{i}\gamma)$ $\leq\epsilon$
$\sum_{i=1}^{T}\log(1-\lambda_{i}\gamma)$ $\leq\log\epsilon$
$\sum_{i=1}^{T}(-\lambda_{i}\gamma)$ $\lesssim\log\epsilon~{}~{}~{}~{}(\log(1-x)\approx-x\text{~{}for small~{}}x)$
$T\times\left(\frac{1}{T}\sum_{i=1}^{T}\lambda_{i}\right)\times\gamma$ $\gtrsim\log\frac{1}{\epsilon}$
Re-arranging this equation gives us the desired result, where $\lambda_{avg}=(\frac{1}{T}\sum_{i}\lambda_{i})$. ∎

### v3

@@ 5.3 Cumulative Weight Decay in Large-Scale Training
Figure 5 depicts the cumulative weight decay $w_{t_{1}}^{t_{2}}$ for three different models: The 124M parameter model from Section 4.2, OLMo 7B [30], and Llama 3 405B [22] (where we assume the model trained with a weight decay of 0.1). The figure depicts the evolution of the term $w_{t_{1}}^{t_{2}}$ for ten fixed choices of $t_{1}$, one at each decile of the training run. Under the assumption that example forgetting occurs at least as fast as the cumulative weight decay, we can interpret the curves in Figure 5 as forgetting curves that indicate to what degree the data seen at different stages is forgotten as we continue training. For the 124M model depicted in Figure 5(a), even the initialization is not completely decayed towards the end (the blue curve is still significantly larger than zero). For OLMo-7B, the gradient weights of the first 40% of the training data decay to zero until the end of training, meaning that this data is likely forgotten (Figure 5(b)). Llama 3 405B also experiences significant decay of the early gradients (Figure 5(c)).
By integrating $w_{t_{1}}^{T}$, the cumulative weight decay allows us to approximate the composition of the final model weights in terms of the gradient updates from different deciles of training. This is depicted in the bottom row of Figure 5.
@@ 6 Discussion
In this work, we have shown that the impact of data contamination depends on the joint scaling of model, data and contamination — an aspect that has largely been overlooked in the literature [71, 57, 40]. Moreover, the impact of contamination can be forgotten as the scale of the data increases, and this is likely the case in many realistic training runs [30, 25].
We have studied contamination with a focus on the leakage of benchmark questions into the training data. This means our work might be more informative than other works that study contamination in different contexts. At the same time, one has to be careful when extrapolating our results, especially to a privacy setup [12, 39]. This is because empirical forgetting might behave differently for random strings or otherwise uniquely identifiable information [13].
The “contamination problem” as studied in this paper has interesting connections to many areas of machine learning, including privacy [29, 39], data attribution [41], and generalization [9, 33, 53]. The connection to data attribution is especially relevant because we show that the presence or absence of a datapoint in the training data can be irrelevant for the model behavior on that same datapoint.
@@ A.3 Filtering Near-Duplicate Benchmark Questions
(a) HellaSwag
(b) WinoGrande
(c) ARC-Easy
(d) PiQA
Figure 7: Contamination and forgetting in OLMo-1B. We contaminate the OLMo-1B checkpoint at gradient step 369,000 four times with different benchmarks. This causes an average accuracy increase of 15 percentage points. We then continue pre-training for 1% of the remaining training time, leading to a reduction of 96% of the accuracy increase due to contamination. In this figure, different colours correspond to different benchmarks, and the grey line depicts the clean accuracy without contamination. Mean and bootstrapped 90% confidence intervals.
(a) Forgetting on Novel Data
(b) After 3 Epochs
(c) After 7 Epochs
Figure 8: Forgetting without weight decay. We perform the forgetting experiment from Section 4.2 without weight decay (instead of the default weight decay of 0.1). This figure depicts the same quantities as Figure 2 in the main paper. We see that there is significant forgetting without weight decay. At the same time, forgetting without weight decay occurs somewhat slower, especially for 144 times repeated contamination. In particular, the accuracy gaps depicted in (b) are slightly larger than the corresponding accuracy gaps in Figure 2, and the difference in (c) is statistically significant (compare Figure 2).
A.4 Reproducibility
The results in this paper were obtained using the OLMo codebase, available at https://github.com/allenai/OLMo, and the llm.c codebase, available at https://github.com/karpathy/llm.c. Our code is fully reproducible, including the random positions at which benchmark questions were inserted into the training data. We trained on the 100BT split of the FineWeb-Edu dataset, available at https://huggingface.co/datasets/HuggingFaceFW/fineweb-edu. Code for this paper is available at https://github.com/tml-tuebingen/forgetting-contamination/. Model training relied on Pytorch [60] and was performed on 8xA100 nodes for all experiments except the continual pre-training of OLMo-7B that ran for a total of 3 weeks on an 8xH100 node.
A.5 Proof of Proposition 1
Fix $\epsilon>0$. Let $T=t_{2}-t_{1}$ be the number of gradient steps that have passed since step $t_{1}$. Let $\gamma_{\text{avg}}=\frac{1}{T}\sum_{t_{1}+1}^{t_{2}}\gamma_{i}$ be the average learning rate between gradient steps $t_{1}$ and $t_{2}$. If $T$ is sufficiently large, that is
$T\geq\frac{\log(1/\epsilon)}{\lambda\gamma_{\text{avg}}}$ (5)
then
$w_{t_{1}}^{t_{2}}\leq\epsilon.$ (6)
@@ A.5 Proof of Proposition 1
Without loss of generality, let $t_{1}=0$ and $T=t_{2}$. We have from equation (4):
$w^{T}_{0}=\prod_{i=1}^{T}(1-\gamma_{i}\lambda).$
We want to guarantee that
$\prod_{i=1}^{T}(1-\lambda_{i}\gamma)\leq\epsilon.$ (7)
First, apply the logarithm
$\sum_{i=1}^{T}\log(1-\gamma_{i}\lambda)\leq\log\epsilon.$ (8)
Now, for $10<x<$ 1 it holds that $\log(1-x)\leq-x$. This means that the inequality (7) is satisfied if
$\lambda\sum_{i=1}^{T}\gamma_{i}\geq-\log\epsilon.$ (9)
Denoting $\gamma_{\text{avg}}=\frac{1}{T}\sum_{i=1}^{T}\gamma_{i}$, this is the same as
$T\geq\frac{\log(1/\epsilon)}{\lambda\gamma_{\text{avg}}}.$